Eight statistics student are given the “Hayashi Test of Extreme Good Looks” (HTEGL). Their scores are the following: 14, 8, 6, 5, 13, 10, 10, 6.

The eight statistics come back 20 years later and have HTEGL scores of 13, 7, 5, 4, 12, 9, 6, 8. Conduct a t-test to see if they are significantly better looking.
Answer #1

Ans:

d=earlier-after 20 years

H_0:\mu _d\geq 0

H_a:\mu _d<0

earlier after 20 d
1 14 13 1
2 8 7 1
3 6 5 1
4 5 4 1
5 13 12 1
6 10 9 1
7 10 6 4
8 6 8 -2
d-bar 1
sd 1.604

Test statistic:

t=(1-0)/(1.604/SQRT(8))

t=1.764

df=8-1=7

p-value=tdist(1.764,7,1)=0.0605

Fail to reject the null hypothesis.

There is not sufficient evidence to conclude that they are significantly better looking after 20 years.